# Covariance scanning for adaptively optimal change point detection in high-dimensional linear models

### v2

@@ B.2.1 Proof of Theorem 3
$\mathsf{P}\left\{|\mathbf{X}^{\top}\widetilde{\mathbf{Y}}(k)-\bar{\mathbf{f}}_{k}|_{\infty}\geq\bar{C}_{0}\|\bm{\Sigma}\|^{1/2}\Psi z\right\}\leq\bar{C}_{1}p\exp(-z^{2})$
@@
where $(\bm{\Sigma}^{1/2})^{\dagger}\bm{\Sigma}(\bm{\Sigma}^{1/2})^{\dagger}$ is a diagonal matrix with $m$ ones and $p-m$ zeros on its diagonal. By Lemma B.5 and $\|\bm{\Sigma}^{1/2}\|=\mu_{1}^{1/2}=\|\bm{\Sigma}\|^{1/2}$, we have for any $0\leq s<e\leq n$ with $e-s\geq\epsilon_{n}$,
$\mathsf{P}\left\{\frac{1}{\sqrt{e-s}}\left|\sum_{t=s+1}^{e}\mathbf{x}_{t}\varepsilon_{t}\right|_{\infty}\geq C\|\bm{\Sigma}\|^{1/2}\sigma z\right\}$
$=$ $\,\mathsf{P}\left\{\frac{1}{\sqrt{e-s}}\max_{i\in[p]}\left|\sum_{t=s+1}^{e}(\bm{\Sigma}^{1/2}\mathbf{e}_{i})^{\top}\mathbf{x}^{\circ}_{t}\varepsilon_{t}\right|\geq C\|\bm{\Sigma}\|^{1/2}\sigma z\right\}\leq 2p\exp(-z^{2}),$ (B.5)
@@
$\mathsf{P}\left\{\frac{1}{\sqrt{e-s}}\left|\sum_{t=s+1}^{e}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\beta}_{j}\right|_{\infty}\geq C\|\bm{\Sigma}\|^{1/2}\Psi z\right\}$
$=$ $\,\mathsf{P}\left\{\frac{1}{\sqrt{e-s}}\max_{i\in[p]}\left|\sum_{t=s+1}^{e}(\bm{\Sigma}^{1/2}\mathbf{e}_{i})^{\top}\left[\mathbf{x}^{\circ}_{t}(\mathbf{x}^{\circ}_{t})^{\top}-(\bm{\Sigma}^{1/2})^{\dagger}\bm{\Sigma}(\bm{\Sigma}^{1/2})^{\dagger}\right]\bm{\Sigma}^{1/2}\bm{\beta}_{j}\right|\geq C\|\bm{\Sigma}\|^{1/2}\Psi z\right\}$
@@
$\mathsf{P}\left\{|\mathbf{X}^{\top}\widetilde{\mathbf{Y}}(k)-\bar{\mathbf{f}}_{k}|_{\infty}\geq\bar{C}_{0}\|\bm{\Sigma}\|^{1/2}\Psi z\right\}$
@@
$\qquad\qquad\qquad\qquad\qquad\left.-\frac{1}{k}\sum_{t=1}^{k}\mathbf{x}_{t}\varepsilon_{t}+\frac{1}{n-k}\sum_{t=k+1}^{n}\mathbf{x}_{t}\varepsilon_{t}\Biggr{|}_{\infty}\geq\bar{C}_{0}\|\bm{\Sigma}\|^{1/2}\Psi z\right\}$
$\leq$ $\,\mathsf{P}\left\{\frac{1}{\sqrt{n-k}}\left|\sum_{t=k+1}^{n}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\beta}_{1}\right|_{\infty}\geq\frac{\bar{C}_{0}}{5}\|\bm{\Sigma}\|^{1/2}\Psi z\right\}$
$\,+\mathsf{P}\left\{\frac{1}{\sqrt{n-k}}\left|\sum_{t=k+1}^{\theta}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\delta}\right|_{\infty}\geq\frac{\bar{C}_{0}}{5}\|\bm{\Sigma}\|^{1/2}\Psi z\right\}$
$\,+\mathsf{P}\left\{\frac{1}{\sqrt{k}}\left|\sum_{t=1}^{k}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\beta}_{0}\right|_{\infty}\geq\frac{\bar{C}_{0}}{5}\|\bm{\Sigma}\|^{1/2}\Psi z\right\}$
$\,+\mathsf{P}\left\{\frac{1}{\sqrt{k}}\left|\sum_{t=1}^{k}\mathbf{x}_{t}\varepsilon_{t}^{\top}\right|_{\infty}\geq\frac{\bar{C}_{0}}{5}\|\bm{\Sigma}\|^{1/2}\sigma z\right\}$
$\,+\mathsf{P}\left\{\frac{1}{\sqrt{n-k}}\left|\sum_{t=k+1}^{n}\mathbf{x}_{t}\varepsilon_{t}^{\top}\right|_{\infty}\geq\frac{\bar{C}_{0}}{5}\|\bm{\Sigma}\|^{1/2}\sigma z\right\}.$ (B.7)
@@
$\mathsf{P}\left\{\frac{1}{\sqrt{n-k}}\left|\sum_{t=k+1}^{\theta}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\delta}\right|_{\infty}\geq C\|\bm{\Sigma}\|^{1/2}\Psi z\right\}$
=\;= $\mathsf{P}\left\{\frac{1}{\sqrt{\theta-k}}\left|\sum_{t=k+1}^{\theta}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\delta}\right|_{\infty}\geq C\|\bm{\Sigma}\|^{1/2}\Psi z\frac{\sqrt{n-k}}{\sqrt{\theta-k}}\right\}$
[…]
@@ B.2.3 Proof of Theorem 5
The proof follows closely the steps in the proofs of Theorems 1 and 2 of Kovács et al., (2024). For notational convenience, we assume that there is no rounding in determining the dyadic search locations and the probe points. Throughout, we omit the subscript ‘Q’ where there is no confusion.
@@
$\leq 2\log_{2}(n)\cdot C_{1}\exp(-2\log\log(n))\leq\frac{2C_{1}}{\log(2)}\log^{-1}(n),$
provided that $\sqrt{p\log\log(n)}\geq 2$. That is, with probability greater than $1-\tfrac{2C_{1}}{\log(2)}\log^{-1}(n)$, Algorithm 1 returns $(\widehat{\theta},\widehat{q})=(n,0)$.
@@
$\mathsf{P}(\mathcal{A}_{1})\geq 1-(4\log^{-1}(2)+2)C_{1}\log^{-1}(n).$
@@
$\quad\left.\text{such that $t,u$ lie on the same side of $\theta$ with $|t-\theta|<|u-\theta|$},\right.$
@@
It can be shown that $\mathcal{P}$ is a subset of a deterministic set, say $\mathcal{Q}$, which depends only on $f_{k}$ defined in (7), and further that $|\mathcal{Q}|\leq 16\log^{2}(n)$ (see page 52 of Kovács et al.,, 2024). The claim on the rate of estimation follows if $\mathsf{P}(\mathcal{A}_{2}^{c})\lesssim\log^{-1}(n)$, where
@@
[…]
@@
$\mathsf{P}(\mathcal{A}_{2})\geq 1-\mathsf{P}(\mathcal{A}_{1}\cap\mathcal{A}_{2}^{c})-\mathsf{P}(\mathcal{A}_{1}^{c})\geq 1-80C_{1}\log^{-1}(n).$
@@ B.2.4 Proof of Proposition 6
$\mathcal{B}_{1}(\mathbf{a})$ :=$:=\left\{\max_{\begin{subarray}{c}0\leq s<e\leq n\end{subarray}}\frac{1}{\sqrt{\max(e-s,\bar{C}_{1}\log(p\vee n))}}\left|\sum_{t=s+1}^{e}\mathbf{a}^{\top}\mathbf{x}_{t}\varepsilon_{t}\right|\leq\bar{C}_{2}|\bm{\Sigma}^{1/2}\mathbf{a}|_{2}\sigma\sqrt{\log(p\vee n)}\right\},$
$\mathcal{B}_{2}(\mathbf{a},\mathbf{b})$ :=$:=\Biggl{\{}\max_{\begin{subarray}{c}0\leq s<e\leq n\end{subarray}}\frac{1}{\sqrt{\max(e-s,\bar{C}_{1}\log(p\vee n))}}\left|\sum_{t=s+1}^{e}\mathbf{a}^{\top}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\mathbf{b}\right|$
@@
$\mathcal{B}_{3}(\mathbf{a})$ :=$:=\left\{\frac{\underline{\sigma}}{2}|\mathbf{a}|_{2}^{2}-\frac{C_{\text{RE}}\log(p)}{\underline{\sigma}n}|\mathbf{a}|_{1}^{2}\leq\frac{1}{n}\sum_{t\in[n]}\mathbf{a}^{\top}\mathbf{x}_{t}\mathbf{x}_{t}^{\top}\mathbf{a}\leq\frac{3\|\bm{\Sigma}\|}{2}|\mathbf{a}|_{2}^{2}+\frac{C_{\text{RE}}\log(p)}{\underline{\sigma}n}|\mathbf{a}|_{1}^{2}\right\},$
@@
where the first inequality follows from the $\varepsilon$-net argument (see e.g. Lemma F.2 of Basu and Michailidis,, 2015), and the second inequality holds with the choice
@@
which is an integer under the condition on $n$. This, in combination of Lemma 12 of Loh and Wainwright, (2012), gives
@@
When either (i) or (ii) holds, by the arguments employed in the proofs of Theorems 3 and 5, we have $|\widehat{\theta}-\theta|\leq\eta\Delta$ for some $\eta\in(0,1)$ that depends on $\bar{c}_{0}$ or $c_{0}$, with large probability. Then, evoking Proposition S3.6 in the supplement of Cho et al., (2025), we have
$\left|\widehat{\bm{\delta}}-\bm{\delta}(\widehat{\theta})\right|_{2}\lesssim\frac{\|\bm{\Sigma}\|^{1/2}\Psi\sqrt{\mathfrak{s}_{\delta}\log(p\vee n)}}{\underline{\sigma}\sqrt{\Delta}}\text{ \ and \ }$
$\left|\widehat{\bm{\delta}}-\bm{\delta}(\widehat{\theta})\right|_{1}\lesssim\frac{\|\bm{\Sigma}\|^{1/2}\Psi\mathfrak{s}_{\delta}\sqrt{\log(p\vee n)}}{\underline{\sigma}\sqrt{\Delta}},$
where $\bm{\delta}(\widehat{\theta})=\bm{\Sigma}^{-1}(\bm{\gamma}_{\widehat{\theta},n}-\bm{\gamma}_{0,\widehat{\theta}})$, deterministically on the event $\mathcal{B}$.
@@
$\leq\frac{\bar{c}_{1}|\bm{\Sigma}\bm{\delta}|_{\infty}^{-2}\|\bm{\Sigma}\|\Psi^{2}\log(p\log(n))\cdot|\bm{\delta}|_{1}}{\Delta}$
$\leq\frac{\bar{c}_{1}\|\bm{\Sigma}\|\Psi^{2}\log(p\log(n))\cdot\mathfrak{s}_{\delta}}{\underline{\sigma}\sqrt{\Delta}\cdot\sqrt{|\bm{\Sigma}\bm{\delta}|_{\infty}^{2}\Delta}}$
$\lesssim\frac{\|\bm{\Sigma}\|^{1/2}\Psi\mathfrak{s}_{\delta}\sqrt{\log(p\log(n))}}{\underline{\sigma}\sqrt{\Delta}},$
@@
$\left|\bm{\delta}(\widehat{\theta}_{\text{\tiny M}})-\bm{\delta}\right|_{2}$ $\leq\frac{\bar{c}_{1}|\bm{\Sigma}\bm{\delta}|_{\infty}^{-2}\|\bm{\Sigma}\|\Psi^{2}\log(p\log(n))\cdot|\bm{\delta}|_{2}}{\Delta}$
$\lesssim\frac{\|\bm{\Sigma}\|^{1/2}\Psi\sqrt{\mathfrak{s}_{\delta}\log(p\log(n))}}{\underline{\sigma}\sqrt{\Delta}}.$

### v3

@@ B.2.1 Proof of Theorem 3
Recall that $\sigma_{X}^{2}=\max_{i\in[p]}\mathsf{Var}(X_{it})$, where $\mathbf{x}_{t}=(X_{1t},\ldots,X_{pt})^{\top}$.
@@
$\mathsf{P}\left\{|\mathbf{X}^{\top}\widetilde{\mathbf{Y}}(k)-\bar{\mathbf{f}}_{k}|_{\infty}\geq\bar{C}_{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi z\right\}\leq\bar{C}_{1}p\exp(-z^{2})$
@@
where $(\bm{\Sigma}^{1/2})^{\dagger}\bm{\Sigma}(\bm{\Sigma}^{1/2})^{\dagger}$ is a diagonal matrix with $m$ ones and pmp-m zeros on its diagonal. By Lemma B.5 and that $|\bm{\Sigma}^{1/2}\mathbf{e}_{i}|_{2}=\sqrt{\mathsf{Var}(X_{it})}\leq\sigma_{X}$, we have for any $0\leq s<e\leq n$ with $e-s\geq\epsilon_{n}$,
$\mathsf{P}\left\{\frac{1}{\sqrt{e-s}}\left|\sum_{t=s+1}^{e}\mathbf{x}_{t}\varepsilon_{t}\right|_{\infty}\geq C{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\sigma z\right\}$
$=$ $\,\mathsf{P}\left\{\frac{1}{\sqrt{e-s}}\max_{i\in[p]}\left|\sum_{t=s+1}^{e}(\bm{\Sigma}^{1/2}\mathbf{e}_{i})^{\top}\mathbf{x}^{\circ}_{t}\varepsilon_{t}\right|\geq C{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\sigma z\right\}\leq 2p\exp(-z^{2}),$ (B.3)
@@
$\mathsf{P}\left\{\frac{1}{\sqrt{e-s}}\left|\sum_{t=s+1}^{e}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\beta}_{j}\right|_{\infty}\geq C{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi z\right\}$
$=$ $\,\mathsf{P}\left\{\frac{1}{\sqrt{e-s}}\max_{i\in[p]}\left|\sum_{t=s+1}^{e}(\bm{\Sigma}^{1/2}\mathbf{e}_{i})^{\top}\left[\mathbf{x}^{\circ}_{t}(\mathbf{x}^{\circ}_{t})^{\top}-(\bm{\Sigma}^{1/2})^{\dagger}\bm{\Sigma}(\bm{\Sigma}^{1/2})^{\dagger}\right]\bm{\Sigma}^{1/2}\bm{\beta}_{j}\right|\geq C{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi z\right\}$
@@
$\mathsf{P}\left\{|\mathbf{X}^{\top}\widetilde{\mathbf{Y}}(k)-\bar{\mathbf{f}}_{k}|_{\infty}\geq\bar{C}_{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi z\right\}$
@@
$\qquad\qquad\qquad\qquad\qquad\left.-\frac{1}{k}\sum_{t=1}^{k}\mathbf{x}_{t}\varepsilon_{t}+\frac{1}{n-k}\sum_{t=k+1}^{n}\mathbf{x}_{t}\varepsilon_{t}\Biggr|_{\infty}\geq\bar{C}_{0}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi z\right\}$
$\leq$ $\,\mathsf{P}\left\{\frac{1}{\sqrt{n-k}}\left|\sum_{t=k+1}^{n}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\beta}_{1}\right|_{\infty}\geq\frac{\bar{C}_{0}}{5}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi z\right\}$
$\,+\mathsf{P}\left\{\frac{1}{\sqrt{n-k}}\left|\sum_{t=k+1}^{\theta}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\delta}\right|_{\infty}\geq\frac{\bar{C}_{0}}{5}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi z\right\}$
$\,+\mathsf{P}\left\{\frac{1}{\sqrt{k}}\left|\sum_{t=1}^{k}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\beta}_{0}\right|_{\infty}\geq\frac{\bar{C}_{0}}{5}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi z\right\}$
$\,+\mathsf{P}\left\{\frac{1}{\sqrt{k}}\left|\sum_{t=1}^{k}\mathbf{x}_{t}\varepsilon_{t}^{\top}\right|_{\infty}\geq\frac{\bar{C}_{0}}{5}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\sigma z\right\}$
$\,+\mathsf{P}\left\{\frac{1}{\sqrt{n-k}}\left|\sum_{t=k+1}^{n}\mathbf{x}_{t}\varepsilon_{t}^{\top}\right|_{\infty}\geq\frac{\bar{C}_{0}}{5}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\sigma z\right\}.$ (B.5)
@@
$\mathsf{P}\left\{\frac{1}{\sqrt{n-k}}\left|\sum_{t=k+1}^{\theta}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\delta}\right|_{\infty}\geq C{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi z\right\}$
$=\;$ $\mathsf{P}\left\{\frac{1}{\sqrt{\theta-k}}\left|\sum_{t=k+1}^{\theta}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\bm{\delta}\right|_{\infty}\geq C{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi z\frac{\sqrt{n-k}}{\sqrt{\theta-k}}\right\}$
$\leq\;$ $2p\exp\left[-\min\left(z^{2}\frac{n-k}{\theta-k},z\sqrt{n-k}\right)\right]\leq 2p\exp(-z^{2}),$
@@ B.2.3 Proof of Theorem 5
The proof follows closely the steps in the proofs of Theorems 1 and 2 of Kovács et al. (2024). For notational convenience, we assume that there is no rounding in determining the dyadic search locations and the probe points. Throughout, we omit the subscript ‘Q’ where there is no confusion.
@@
$\leq 2\log_{2}(n)\cdot C_{1}\exp(-2\log\log(n))\leq\frac{2C_{1}}{\log(2)}\bigl(\log(n)\bigr)^{-1},$
provided that $\sqrt{p\log\log(n)}\geq 2$. That is, with probability greater than $1-\tfrac{2C_{1}}{\log(2)}\bigl(\log(n)\bigr)^{-1}$, Algorithm 1 returns $(\widehat{\theta},\widehat{q})=(n,0)$.
@@
$\mathsf{P}(\mathcal{A}_{1})\geq 1-(4\log^{-1}(2)+2)C_{1}\bigl(\log(n)\bigr)^{-1}.$
@@
Based on Step 1, provided that $\mathcal{A}_{1}$ holds, it is sufficient to consider the case where we have $(s,e)=(k^{*}/2,2k^{*})$ at the start of the OS algorithm. Besides, it is enough to consider the pairs of the probe points $(t,u)$ considered by the OS algorithm which lie on the same side of $\theta$, since otherwise no matter which side is dropped off, no mistake occurs. Thus we define
@@
It can be shown that $\mathcal{P}$ is a subset of a deterministic set, say $\mathcal{Q}$, which depends only on $f_{k}$ defined in (7), and further that $|\mathcal{Q}|\leq 16\log^{2}(n)$ (see page 52 of Kovács et al., 2024). The claim on the rate of estimation follows if $\mathsf{P}(\mathcal{A}_{2}^{c})\lesssim\bigl(\log(n)\bigr)^{-1}$, where
@@
$\leq\,$ $4C_{1}\exp\left\{-\left[\frac{|\bm{\Sigma}\bm{\delta}|_{2}^{2}(u-t)\theta^{2}}{2C_{0}tu\|\bm{\Sigma}\|\Psi^{2}\sqrt{p}}\right]^{2}\right\}$
@@
$\mathsf{P}(\mathcal{A}_{2})\geq 1-\mathsf{P}(\mathcal{A}_{1}\cap\mathcal{A}_{2}^{c})-\mathsf{P}(\mathcal{A}_{1}^{c})\geq 1-80C_{1}\bigl(\log(n)\bigr)^{-1}.$
@@ B.2.4 Proof of Proposition 6
$\mathcal{B}_{1}(\mathbf{a})$ $:=\left\{\max_{\begin{subarray}{c}0\leq s<e\leq n\end{subarray}}\frac{1}{\sqrt{\max(e-s,\bar{C}_{1}\log(p\vee n))}}\left|\sum_{t=s+1}^{e}\mathbf{a}^{\top}\mathbf{x}_{t}\varepsilon_{t}\right|\leq\bar{C}_{2}|\bm{\Sigma}^{1/2}\mathbf{a}|_{2}\sigma\sqrt{\log(p\vee n)}\right\},$
$\mathcal{B}_{2}(\mathbf{a},\mathbf{b})$ $:=\Biggl\{\max_{\begin{subarray}{c}0\leq s<e\leq n\end{subarray}}\frac{1}{\sqrt{\max(e-s,\bar{C}_{1}\log(p\vee n))}}\left|\sum_{t=s+1}^{e}\mathbf{a}^{\top}(\mathbf{x}_{t}\mathbf{x}_{t}^{\top}-\bm{\Sigma})\mathbf{b}\right|$
@@
$\mathcal{B}_{3}(\mathbf{a})$ $:=\left\{\frac{\underline{\sigma}}{2}|\mathbf{a}|_{2}^{2}-\frac{C_{\text{RE}}\log(p)}{\underline{\sigma}n}|\mathbf{a}|_{1}^{2}\leq\frac{1}{n}\sum_{t\in[n]}\mathbf{a}^{\top}\mathbf{x}_{t}\mathbf{x}_{t}^{\top}\mathbf{a}\leq\frac{3\|\bm{\Sigma}\|}{2}|\mathbf{a}|_{2}^{2}+\frac{C_{\text{RE}}\log(p)}{\underline{\sigma}n}|\mathbf{a}|_{1}^{2}\right\},$
@@
where the first inequality follows from the $\varepsilon$-net argument (see e.g. Lemma F.2 of Basu and Michailidis, 2015), and the second inequality holds with the choice
@@
which is an integer under the condition on $n$. This, in combination of Lemma 12 of Loh and Wainwright (2012), gives
@@
When either (i) or (ii) holds, by the arguments employed in the proofs of Theorems 3 and 5, we have $|\widehat{\theta}-\theta|\leq\eta\Delta$ for some $\eta\in(0,1)$ that depends on $\bar{c}_{0}$ or $c_{0}$, with large probability. Then, evoking Proposition S3.6 in the supplement of Cho et al. (2025), we have
$\left|\widehat{\bm{\delta}}-\bm{\delta}(\widehat{\theta})\right|_{2}\lesssim\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi\sqrt{\mathfrak{s}_{\delta}\log(p\vee n)}}{\underline{\sigma}\sqrt{\Delta}}\text{ \ and \ }$
$\left|\widehat{\bm{\delta}}-\bm{\delta}(\widehat{\theta})\right|_{1}\lesssim\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi\mathfrak{s}_{\delta}\sqrt{\log(p\vee n)}}{\underline{\sigma}\sqrt{\Delta}},$
[…]
@@
$\leq\frac{\bar{c}_{1}|\bm{\Sigma}\bm{\delta}|_{\infty}^{-2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}^{2}}\Psi^{2}\log(p\log(n))\cdot|\bm{\delta}|_{1}}{\Delta}$
$\leq\frac{\bar{c}_{1}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}^{2}}\Psi^{2}\log(p\log(n))\cdot\mathfrak{s}_{\delta}}{\underline{\sigma}\sqrt{\Delta}\cdot\sqrt{|\bm{\Sigma}\bm{\delta}|_{\infty}^{2}\Delta}}$
$\lesssim\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi\mathfrak{s}_{\delta}\sqrt{\log(p\log(n))}}{\underline{\sigma}\sqrt{\Delta}},$
@@
$\left|\bm{\delta}(\widehat{\theta}_{\text{\tiny M}})-\bm{\delta}\right|_{2}$ $\leq\frac{\bar{c}_{1}|\bm{\Sigma}\bm{\delta}|_{\infty}^{-2}{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}^{2}}\Psi^{2}\log(p\log(n))\cdot|\bm{\delta}|_{2}}{\Delta}$
$\lesssim\frac{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\sigma_{X}}\Psi\sqrt{\mathfrak{s}_{\delta}\log(p\log(n))}}{\underline{\sigma}\sqrt{\Delta}}.$